1.1 The operations manager of a sugar mill in Chesterville wants to estimate the average size of an order received. An order is measured in the number of pallets shipped from overseas. A random sample of 84 orders from customers had a sample mean value of 121.6 pallets. Assume that the population standard deviation is 20 pallets and that order size is normally distributed. Estimate, with 92% confidence, the mean size of orders received from all the mill’s customers. 


1
Expert's answer
2021-09-01T10:18:05-0400

The critical value for "\alpha=0.08" is "z_c=z_{1-\alpha/2}=1.7507."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-z_c\times\dfrac{\sigma}{
\sqrt{n}}, \bar{X}+z_c\times\dfrac{\sigma}{
\sqrt{n}})"

"=(121.6-1.7507\times\dfrac{20}{
\sqrt{84}}, 121.6+1.7507\times\dfrac{20}{
\sqrt{84}})"

"=(117.78,125.42)"

Therefore, based on the data provided, the 92 % confidence interval for the population mean is "117.78<\mu<125.42," which indicates that we are 92% confident that the true population mean "\mu" is contained by the interval "(117.78,125.42)."